# TriVLA: A Triple-System-Based Unified Vision-Language-Action Model for General Robot Control

### v2

@@ Abstract
Recent advancements in vision-language models (VLMs) for common-sense reasoning have led to the development of vision-language-action (VLA) models, enabling robots to perform generalized manipulation. Although existing autoregressive VLA methods design a specific architecture like dual-system to leverage large-scale pretrained knowledge, they tend to capture static information, often neglecting the dynamic aspects vital for embodied tasks. To this end, we propose TriVLA, a unified Vision-Language-Action model with a triple-system architecture for general robot control. The vision-language module (System 2) interprets the environment through vision and language instructions. The dynamics perception module (System 3) inherently produces visual representations that encompass both current static information and predicted future dynamics, thereby providing valuable guidance for policy learning. TriVLA utilizes pre-trained VLM model and fine-tunes pre-trained video foundation model on robot datasets along with internet human manipulation data. The subsequent policy learning module (System 1) generates fluid motor actions in real time. Experimental evaluation demonstrates that TriVLA operates at approximately 36 Hz and surpasses state-of-the-art imitation learning baselines on standard simulation benchmarks as well as challenging real-world manipulation tasks. [^1] Project in https://zhenyangliu.github.io/TriVLA/
@@ 1 Introduction
Figure 1: Our TriVLA is a Vision-Language-Action (VLA) system with a triple-system architecture. The TriVLA converts image observations and language instructions into token sequences, which are processed by the Vision-Language Model (VLM) for reasoning using common knowledge, and the Stable Video Diffusion Model (VDM) as the world model for both current and future predictions. The outputs of the VLM and VDM, along with robot state and action encodings, are fed into the policy learning module to generate motor actions. The TriVLA can be used directly to perform tasks based on prompts or fine-tuned with high-quality data to support complex multi-stage tasks.
Recent vision-language models (VLMs) Liu et al. (2024c); Alayrac et al. (2022); Li et al. (2023a); Zhang et al. (2023); Bai et al. (2023); Gao* et al. (2023); Zhang et al. (2024b; a) have exhibited notable abilities in instruction following and commonsense reasoning. These advances are largely attributed to pretraining on extensive image-text datasets sourced from the internet. Building upon these advancements, recent works have proposed specialized architectures—including dual-system approaches—to augment VLMs into vision-language-action (VLA) models. These models facilitate action plan generationAhn et al. (2022); Driess et al. (2023); Huang et al. (2023); Belkhale et al. (2024) and SE(3) pose estimationBrohan et al. (2023); Kim et al. (2024); Li et al. (2024a). VLA models empower robots to interpret visual inputs and linguistic commands, thereby generating generalizable control actions. Existing VLMs for vision and language Li et al. (2024b) serve as foundational components for training robotic policies capable of generalizing to novel objects, environments, and tasks beyond the training distribution.
Although VLM-based VLA approaches Intelligence et al. (2025); Black et al. (2024); Brohan et al. (2023); Kim et al. (2024); Pertsch et al. (2025) have achieved success in embodied tasks, they predominantly focus on static information. This limitation arises because these methods typically utilize only one or two current sampled images, thereby underutilizing the dynamic information inherent in sequential demonstrations. It is essential to leverage the intrinsic strengths of VLMs to develop VLA models capable of robust manipulation within dynamic environments. Recent advances in video diffusion models (VDMs) Blattmann et al. (2023a); Hong et al. (2022); Yang et al. (2024); Brooks et al. (2024) have demonstrated remarkable performance in video generation tasks. Unlike methods pretrained on single images or image pairs, VDMs model entire video sequences and predict future frames guided by current observations and instructions, thereby exhibiting strong comprehension of physical dynamics.
Inspired by the strong prediction and dynamic perception capabilities of VDMs, we introduce TriVLA, a unified Vision-Language-Action model to integrate the physical dynamics into general policy learning. As shown in Figure 2, TriVLA adopts a triple-system compositional architecture on the basis of the existing dual-system structure Bjorck et al. (2025); Shi et al. (2025). The System 2 vision-language module is a pre-trained Vision-Language Model (VLM) that processes the robot’s visual perception and language instruction to interpret the environment and understand the task goal. The System 3 Dynamics Perception Module is a versatile video diffusion model fine-tuned on large-scale human and robotic manipulation data collected from the internet Khazatsky et al. (2024); Jin et al. (2024); Lu et al. (2024). This module focuses on developing a controllable video generation model to enhance predictive accuracy within the manipulation domain. Consequently, the downstream policy can implicitly acquire the inverse dynamics model Min et al. (2023); Tian et al. (2024) by monitoring the robot’s motions embedded in the predictive representation. This approach enables transfer of the video prediction model’s generalization capabilities to robotic policies. Only a limited number of demonstrations are required to align the robot’s action space with the visual domain. Subsequently, the System 1 policy learning module, trained via action flow-matching, cross-attends to output tokens from Systems 2 and 3. It utilizes embodiment-specific encoders and decoders to manage variable state and action dimensions during motion generation. Motivated by recent progress in robot learning, the System 1 policy learning module predicts action sequences (action chunking) instead of single actions at each timestep.
Figure 2: Comparison between previous dual-system architectures and our triple-system approach. Our TriVLA employs a unified triple-system compositional architecture that integrates world knowledge (System 2) and the world model (System 3), both critical for general policy learning. Prior dual-system methods typically addressed only one component, failing to unify both.
The primary contribution of this paper is the introduction of a triple-system compositional architecture. This innovative framework integrates Vision-Language Models (VLMs) and Video Diffusion Models (VDMs) to support both high-level reasoning based on common knowledge and dynamic predictive representation through a world model. As illustrated in Figure 1, the proposed unified framework allows robots to handle significantly more complex prompts and endows them with long-horizon task execution capabilities. Although individual components of the triple-system, including the VLM and VDM modules, have been explored in previous studies, their integration is novel and facilitates unique functionalities.
@@
Experimental results demonstrate that the proposed TriVLA consistently outperforms baseline algorithms—including API-based VLMs and flat VLA policies—across three simulated Mees et al. (2022); Liu et al. (2024a); Yu et al. (2020) and three real-world environments. This highlights its effectiveness in aligning with human intent and achieving long-horizon task success. Notably, TriVLA attains improvements of 0.21, 0.11, and 0.13 on the Calvin ABC→D, LIBERO, and MetaWorld benchmarks, respectively, compared to prior state-of-the-art methods.
@@
  • A Unified Vision-Language-Action Framework: We propose a unified Vision-Language-Action model to integrate the world knowledge and world model for general policy learning across multiple robot embodiments.
  • Triple-System Compositional Architecture: The proposed TriVLA model designs a novel triple-system compositional architecture that possesses both high-level reasoning and dynamic predictive representation, enables a robot to process much more complex prompts and long-horizon manipulation tasks.
  • State-of-the-art performance: TriVLA outperforms other baseline algorithms across simulated and real-world settings, including new combinations of skills seen during training, in the context of scenario. This demonstrates the effectiveness in both alignment with human intent and long-horizon task success.
@@ 2 Related Work
Previous works Ahn et al. (2022); Driess et al. (2023); Huang et al. (2023; 2024b) facilitate robotic understanding of language instructions and visual inputs to autonomously generate task plans. Concurrently, vision-language-action (VLA) models exploit the innate reasoning capacity of VLMs to predict low-level SE(3) poses. In particular, RT2 Brohan et al. (2023) discretizes 7-DoF actions into bins for autoregressive pose prediction. Extending this approach, ManipLLM Li et al. (2024a) integrates affordance priors via chain-of-thought reasoning, whereas OpenVLA Kim et al. (2024) conducts large-scale pretraining on the Open X-Embodiment dataset O’Neill et al. (2023), thereby improving generalization. FAST Pertsch et al. (2025) utilizes the discrete cosine transform to facilitate efficient and scalable training of autoregressive VLA models. The dual-system architectures of GR00T N1 Bjorck et al. (2025) and Hi Robot Shi et al. (2025), motivated by human cognition models, empower robots to reason about unfamiliar scenarios, robustly manage real-world variability, and accelerate learning of new tasks. Several VLA methods Liu et al. (2024d); Huang et al. (2024a); Li et al. (2023b); Wu et al. (2023) enable continuous action prediction by integrating policy heads like MLPs or LSTMs Graves & Graves (2012), and applying regression loss within imitation learning frameworks. Nevertheless, these approaches predominantly capture static information and insufficiently leverage dynamics present in sequential demonstrations, largely due to reliance on only one or two current sampled images. Therefore, exploiting the intrinsic strengths of VLMs to construct VLA models capable of robust manipulation in dynamic environments is imperative. Our proposed TriVLA employs a triple-system architecture to incorporate physical dynamics into policy learning.
@@
Prior studies have investigated leveraging future prediction to improve policy learning Bharadhwaj et al. (2024); Chen et al. (2024); Ye et al. (2024); Guo et al. (2024). SuSIE Black et al. (2023) bases its control policy on a predicted future keyframe produced by InstructPix2Pix Brooks et al. (2023), whereas UniPi Du et al. (2024) models inverse dynamics across two generated frames. These approaches rely on a single-step future prediction to guide action selection, failing to fully capture the complexity of physical dynamics. Furthermore, denoising the final predicted image is time-consuming and results in reduced control frequency. GR-1 Wu et al. (2023) generates future frames and actions in an autoregressive manner. However, it produces only one image per forward pass, and its prediction quality trails behind diffusion-based techniques. Seer Tian et al. (2024) proposes an end-to-end framework that predicts actions by applying inverse dynamics models conditioned on forecasted visual states. VPP Hu et al. (2024) employs representations fine-tuned from video foundation models to build a generalist robotic policy. In contrast, our TriVLA incorporates a fine-tuned video foundation model as an additional System 3 within the dual-system VLM-based VLA architecture. This enables prediction of sequential future frames while maintaining high-level reasoning, thereby more effectively guiding policy learning.
Figure 3: The pipeline of our TriVLA. Our TriVLA adopts a triple-system compositional architecture on the basis of the existing dual-system structure. The System 2 vision-language module employs a pre-trained Eagle-2 Vision-Language Model (VLM) to process the robot’s visual inputs and language instructions, enabling environmental interpretation and task goal understanding. The System 3 dynamics perception module uses a general-purpose video diffusion model to capture entire video sequences and predict future frames based on current observations and task instructions. Subsequently, the System 1 policy learning module—trained using action flow-matching—cross-attends to the output tokens from Systems 2 and 3, and employs embodiment-specific encoders and decoders to handle variable state and action dimensions for generating motor actions.
@@ 3 Preliminaries
Vision-language-action model. Robotic manipulation has long been a central focus within the robotics community. Vision-language-action (VLA) models are designed to iteratively determine the robot’s subsequent action—specifically, the end-effector’s pose—based on visual observations and human instructions that define the current task. Recently, high-capacity, pretrained vision-and-language models (VLMs) have demonstrated strong generalization capabilities across diverse language-conditioned manipulation tasks. Among these, recent VLM-based VLA methods adopt a dual-system compositional architecture inspired by human cognitive processing Kahneman (2011). This architecture facilitates higher-level reasoning to interpret complex long-horizon tasks and select appropriate action sequences. At each timestep $t$, The high-level system takes the captured image $\mathbf{o}_{t}$ by cameras from base and wrist mounted cameras and the open-ended instruction $\mathbf{v}_{t}^{\text{in}}$ as input to generate reasoning tokens. The low-level policy uses these tokens, images, and robot states and outputs an action token sequence $\mathbf{v}_{t}^{\text{out}}\in\mathcal{V}^{n}$, where each action token represents a discrete bin of one dimension of the robot action space. The final robot action is extracted from this sequence using a post-processing function $f$, resulting in $\mathbf{a}_{t}=f(\mathbf{v}_{t}^{\text{out}})$.
@@ 4 Our Proposed TriVLA
This section outlines the triple-system learning procedure of TriVLA, as illustrated in Figure 3. Initially, the NVIDIA Eagle-2 VLM Li et al. (2025) is employed as the vision-language backbone to process visual perception and language instructions. Subsequently, a video diffusion model fine-tuned on diverse manipulation datasets is adopted to generate visual representations. TriVLA employs flow matching Lipman et al. (2022) to facilitate action generation via a diffusion transformer (DiT).
4.1 Vision-Language Module (System 2)
@@
To encode vision and language inputs, TriVLA employs the Eagle-2 vision-language model (VLM) pretrained on large-scale internet data. Eagle-2 is fine-tuned from a SmolLM2 large language model (LLM) and a SigLIP-2 image encoder. Images are encoded at a resolution of 224 × 224 pixels, followed by pixel shuffle, producing 64 image token embeddings per frame. These embeddings are subsequently encoded jointly with textual input by the LLM component of the Eagle-2 VLM. The LLM and image encoder are aligned over a wide range of vision-language tasks according to a general training protocol. During policy training, task descriptions in text form, along with potentially multiple images, are input to the VLM using the chat format implemented during vision-language pretraining. Vision-language tokens, denoted as $Q_{vl}$ with dimensions (batch size × sequence length × hidden dimension), are then extracted from the LLM. Empirically, using embeddings from the middle layers of the LLM rather than the final layer yields faster inference and improved downstream policy success rates. In TriVLA, representations are extracted specifically from the 12th LLM layer. To handle varying state dimensions across diverse robots, TriVLA employs an embodiment-specific MLP to project states into a shared embedding space, denoted as the state token $Q_{s}$.
4.2 Dynamics Perception Module (System 3)

### v3

# TriVLA: A Triple-System-Based Unified Vision-Language-Action Model with Episodic World Modeling for General Robot Control

@@ Abstract
Recent advances in vision–language models (VLMs) have enabled robots to follow open-ended instructions and demonstrate impressive commonsense reasoning. However, current vision–language–action (VLA) frameworks primarily rely on static representations and limited temporal context, restricting agents to short-horizon, reactive behaviors and hindering robust generalization in dynamic embodied environments. Inspired by cognitive neuroscience theories of episodic memory, we are, to our knowledge, among the first to introduce a formalized episodic world model in VLA, enabling embodied robots to accumulate, recall, and predict sequential experiences. As an instantiation of this concept, our unified TriVLA realizes the episodic world model through a triple-system architecture: integrating multimodal grounding from a pretrained VLM (System 2) and temporally rich dynamics perception from a video diffusion model (System 3). This enables the agent to accumulate and recall sequential experiences, interpret current contexts, and predict future environmental evolution. Guided by episodic representations that span both the past and anticipated future, the downstream policy (System 1) generates coherent, context-aware action sequences through flow-matching and cross-modal attention mechanisms. Experimental results show that TriVLA operates efficiently at  36 Hz and consistently outperforms baseline models on standard benchmarks and challenging real-world manipulation tasks. It demonstrates strong long-horizon planning and open-ended intent understanding, showcasing the advantages of episodic world model-inspired reasoning for robust, generalizable robot intelligence. Project Page: zhenyangliu.github.io/TriVLA.
Figure 1: TriVLA is a unified Vision-Language-Action framework that adopts a triple-system architecture inspired by the episodic world model. Image and language inputs are processed by a Vision-Language Model for multimodal perception. A Video Diffusion Model provides dynamic world modeling and future prediction. The policy module integrates sequential outputs, robot state, and action history and generates real-time actions for complex manipulation tasks.
@@ 1 Introduction
“Episodic memory is the only memory system that allows mental time travel—backward into the past and forward into the future.”
— Endel Tulving
Building on this cognitive foundation, we advocate that robotic agents, require an internal episodic world model: a representational system that not only recalls past interactions but also anticipates future dynamics, thereby enabling robust generalization in embodied environments.
Decades of cognitive neuroscience provide compelling evidence for this perspective. As first articulated by Tulving Tulving et al. (1972), episodic memory refers to the encoding, storage, and retrieval of experiences within their spatiotemporal context. This unique system empowers humans not only to recollect the past but also to simulate potential futures, thus grounding flexible planning and adaptive decision-making. Converging findings highlight the central roles of the hippocampus and prefrontal cortex in supporting episodic memory, enabling individuals to integrate sensory cues with temporal dynamics to construct predictive internal world models Tulving (2002); Pritzel et al. (2017); Blundell et al. (2016); Lin et al. (2018); Gershman & Daw (2017). Episodic memory thus forms a fundamental component of intelligence, providing both the experiential basis for learning and the representational scaffolding for generalization across novel tasks.
Inspired by these insights, we propose the concept of an episodic world model: a unified framework that integrates multimodal grounding with temporally rich dynamic modeling. Unlike static scene representations, an episodic world model continuously accumulates, recalls, and predicts sequential experiences, equipping artificial agents with a memory system more akin to human intelligence. Recent advances in video diffusion models (VDMs) Blattmann et al. (2023a); Hong et al. (2022); Yang et al. (2024); Brooks et al. (2024) provide a technological foundation for this paradigm, as they capture temporal continuity and physical dynamics across video sequences, naturally aligning with episodic memory principles and enabling richer, context-aware internal representations.
In parallel, vision–language models (VLMs) Liu et al. (2024c); Alayrac et al. (2022); Li et al. (2023a); Zhang et al. (2023); Bai et al. (2023); Gao* et al. (2023); Zhang et al. (2024b; a) have demonstrated impressive progress in instruction following and commonsense reasoning through large-scale pretraining on image–text corpora. Extending these capabilities, dual-system architectures have advanced VLMs into vision–language–action (VLA) models that generate action plans Ahn et al. (2022); Driess et al. (2023); Huang et al. (2023); Belkhale et al. (2024) and estimate SE(3) object poses Brohan et al. (2023); Kim et al. (2024); Li et al. (2024), enabling robots to map multimodal inputs into generalizable control behaviors. As illustrated in Figure 2, current VLM-based VLA systems Intelligence et al. (2025); Black et al. (2024); Brohan et al. (2023); Kim et al. (2024); Pertsch et al. (2025) remain predominantly static: they depend on one or two instantaneous observations, overlooking the sequential and dynamic structures that characterize embodied interaction. As a result, they cannot encode or utilize temporally extended experiences, a capability similar to human episodic memory and crucial for robust performance in dynamic environments.
To bridge this gap, we introduce TriVLA, a unified Vision–Language–Action model that implements the episodic world model through a triple-system compositional architecture. Extending prior dual-system designs Bjorck et al. (2025); Shi et al. (2025), TriVLA explicitly integrates:
  • System 2: Episodic Multimodal Perception, a pretrained VLM that interprets observations and instructions, summarizing task goals and contextual cues.
  • System 3: Episodic Dynamics Perception, a video diffusion model fine-tuned on large-scale human and robotic manipulation datasets Khazatsky et al. (2024); Jin et al. (2024); Lu et al. (2024), which encodes sequences of past states and predicts future scene trajectories, realizing episodic context accumulation.
Together, Systems 2 and 3 jointly compose the episodic world model, fusing descriptive multimodal grounding with predictive temporal modeling. This joint representation empowers Policy Learning (System 1) to flexibly adapt actions based on accumulated experiences and anticipated future dynamics. By continually monitoring motion sequences embedded within the episodic world representation, the downstream policy induces an implicit inverse-dynamics prior Min et al. (2023); Tian et al. (2024). This facilitates the transfer of the generalization capabilities inherent in the episodic world model to the robotic policy, meaning that the robot requires only a limited number of demonstrations to align its action space with the visual domain. During training, System 1 utilizes action flow-matching and cross-attention mechanisms to integrate output tokens from both Systems 2 and 3. It adopts embodiment-specific encoders and decoders to manage variable state and action dimensions during motion generation. In addition, inspired by recent advances in robot learning, System 1 is designed to predict a chunk of actions rather than generating isolated actions at each timestep.
Figure 2: Comparison between dual-system architectures and our episodic world model-guided TriVLA. TriVLA implements the episodic world model using a triple-system architecture. In contrast, previous dual-system methods relied on static representations and limited temporal context, which restricted agents to short-horizon, reactive behaviors in dynamic environments.
Our main contribution is an episodic world model inspired by cognitive neuroscience theories of episodic memory. Building on this, we introduce a novel triple-system architecture within a unified Vision-Language-Action framework. The episodic world model that guides manipulation policy learning with temporally extended, context-aware signals. As illustrated in Figure 1, this unified framework empowers robots to interpret complex prompts, reason over long horizons, and adapt adaptively in open-ended, dynamic scenarios.
@@
Experimental results demonstrate that the proposed TriVLA consistently outperforms baseline algorithms, in both simulated Mees et al. (2022); Liu et al. (2024a); Yu et al. (2020) and real-world environments. This highlights its effectiveness in aligning with human intent and achieving long-horizon task success. Notably, TriVLA attains improvements of 0.21, 0.11, and 0.13 on the Calvin ABC→D, LIBERO, and MetaWorld benchmarks, respectively, compared to prior state-of-the-art methods. In real-world experiments, TriVLA demonstrates strong effectiveness in dexterous hand manipulation tasks, particularly in long-horizon scenarios.
@@
  • Episodic World Model Inspired by Cognitive Neuroscience: We propose an episodic world model for embodied agents, inspired by cognitive neuroscience theories of episodic memory. This model enables robots to accumulate, recall, and predict sequential multimodal experiences and offers a solid foundation for robust, adaptive control.
  • A Unified Triple-System Compositional Architecture: Building on this foundation, we present TriVLA, a triple-system architecture implementing the episodic world model. TriVLA provides high-level reasoning and dynamic prediction. Robots using TriVLA can understand complex prompts and perform long-horizon manipulation.
  • State-of-the-art Performance: TriVLA outperforms other baseline algorithms, including novel skill compositions beyond training combinations. This demonstrates the effectiveness in both alignment with human intent and long-horizon task success.
@@ 2 Related Work
Previous studies Ahn et al. (2022); Driess et al. (2023); Huang et al. (2023; 2024b) have advanced robotic comprehension of language and vision to autonomously generate task plans. Vision-language-action (VLA) models leverage VLMs’ reasoning capabilities for SE(3) pose prediction: RT2 Brohan et al. (2023) binarizes 7-DoF actions for autoregressive prediction; ManipLLM Li et al. (2024) incorporates affordance priors and chain-of-thought reasoning; OpenVLA Kim et al. (2024) pretrains on Open X-Embodiment O’Neill et al. (2023) for improved generalization; and FAST Pertsch et al. (2025) introduces discrete cosine transform for efficient scalability. Cognitively inspired dual-systems like GR00T N1 Bjorck et al. (2025) and Hi Robot Shi et al. (2025) enhance adaptation to novel scenarios and accelerate task learning. Additionally, various VLA approaches Liu et al. (2024d); Huang et al. (2024a); Li et al. (2023b); Wu et al. (2023) achieve continuous action prediction by integrating policy heads (e.g., MLP, LSTM Graves & Graves (2012)) with regression losses in imitation learning. Previous approaches are mostly static. They rely on current observations and ignore sequential dynamics. As a result, these methods cannot encode temporally extended experiences, which are crucial for robust performance in dynamic environments. In contrast, TriVLA uses a unified triple-system architecture. This design enables robots to interpret complex prompts and perform long-horizon manipulation tasks across various scenarios.
@@
Prior studies have investigated leveraging future prediction to improve policy learning Bharadhwaj et al. (2024); Chen et al. (2024); Ye et al. (2024); Guo et al. (2024). SuSIE Black et al. (2023) bases its control policy on a predicted future keyframe produced by InstructPix2Pix Brooks et al. (2023), whereas UniPi Du et al. (2024) models inverse dynamics across two generated frames. These approaches rely on a single-step future prediction to guide action selection, failing to fully capture the complexity of physical dynamics. Furthermore, denoising the final predicted image is time-consuming and results in reduced control frequency. GR-1 Wu et al. (2023) generates future frames and actions in an autoregressive manner. However, it produces only one image per forward pass, and its prediction quality trails behind diffusion-based techniques. Seer Tian et al. (2024) proposes an end-to-end framework that predicts actions by applying inverse dynamics models conditioned on forecasted visual states. VPP Hu et al. (2024) employs representations fine-tuned from video models to build a generalist robotic policy. In contrast, our TriVLA integrates Episodic Multimodal Perception and Episodic Dynamics Perception to provide both high-level reasoning and dynamic predictive representations. This enables prediction of sequential future frames while maintaining high-level reasoning, thereby more effectively guiding policy learning.
@@ 3 Preliminaries
Vision-language-action model. Robotic manipulation is a longstanding core topic in robotics. Vision-language-action (VLA) models infer the robot’s next action—typically the end-effector pose—from visual input and human instructions. Recent advances in large pretrained vision-language models (VLMs) have enabled strong generalization across diverse, language-conditioned manipulation tasks. Most VLM-based VLA approaches adopt a dual-system architecture inspired by human cognition Kahneman (2011), facilitating higher-level reasoning for complex, long-horizon task interpretation and action selection. At each timestep $t$, the high-level system takes the captured image $\mathbf{o}_{t}$ from base and wrist-mounted cameras, along with the open-ended instruction $\mathbf{v}_{t}^{\text{in}}$, as input to generate reasoning tokens. The low-level policy uses these tokens, images, and robot states and outputs an action token sequence $\mathbf{v}_{t}^{\text{out}}\in\mathcal{V}^{n}$, where each action token represents a discrete bin of one dimension of the robot action space. The final robot action is extracted from this sequence using a post-processing function $f$, resulting in $\mathbf{a}_{t}=f(\mathbf{v}_{t}^{\text{out}})$.
@@ 4 Our Proposed TriVLA
TriVLA implements the episodic world model using a triple-system design (Figure 3). It integrates (i) Episodic Multimodal Perception, which uses the Eagle-2 VLM Li et al. (2025) to interpret visual inputs and language instructions, and (ii) Episodic Dynamics Perception, where a video diffusion model fine-tuned on large-scale manipulation datasets predicts future scene trajectories. These modules provide rich episodic context for policy generalization. The policy module leverages this context with flow matching Lipman et al. (2022) and a diffusion transformer (DiT) to generate actions.
Figure 3: The pipeline of TriVLA. TriVLA is a unified Vision-Language-Action framework built on a triple-system paradigm. System 2 employs a pre-trained Eagle-2 VLM for episodic multimodal perception, while System 3 utilizes a general-purpose VDM to model episodic dynamics and sequential changes. Together, these modules form a joint episodic world model with rich, temporally extended representations. System 1 serves as the policy module, applying action flow-matching to integrate all outputs along with robot state and action history.
4.1 Episodic Multimodal Perception (System 2)
@@
To realize this component, TriVLA employs the NVIDIA Eagle-2 vision–language model (VLM) Li et al. (2025), pretrained on large-scale internet corpora to jointly interpret observations and instructions, summarizing task goals and contextual cues. Eagle-2 is built from a SmolLM2 language model and a SigLIP-2 image encoder, aligned through a broad vision–language pretraining protocol. Images are encoded at a resolution of 224 × 224 pixels and then processed via pixel shuffle to produce 64 image token embeddings per frame. These image embeddings are jointly encoded with textual input by the LLM component of Eagle-2. The LLM and image encoder are aligned through a general vision-language pretraining protocol covering diverse tasks. During policy training, task descriptions and multiple images are input to the VLM following the chat format established during pretraining. The resulting vision-language tokens, denoted as $Q_{vl}$ (batch size × sequence length × hidden dimension), are extracted from the LLM. Empirically, extracting embeddings from the 12th layer of the LLM, rather than from the final layer, yields faster inference and higher policy success rates. To handle varying robot state dimensions, TriVLA uses an embodiment-specific MLP to project each robot’s state into a shared embedding space, resulting in a state token $Q_{s}$.
4.2 Episodic Dynamics Perception (System 3)